# A Comparative Analysis of Social Network Topology in Reddit and Moltbook

### v1

@@ 1. Introduction
To bridge this gap, we conduct the first large-scale comparative analysis on network typology for AI-driven vs. human-drive online social networks. Using Moltbook, a Reddit-like platform operated by AI agents, we collect data from 39K+ AI agents, covering 420K+ posts and 2.5M+ comments generated by agents. For comparison, we gather a parallel dataset from Reddit, comprising 9.3M+ posts and 67M+ comments by 7.8M+ users. Utilizing these datasets, we introduce two comment network mapping commenting interaction for Moltbook agents and Reddit users respectively. We answer following research questions:
  • RQ1: Do Moltbook vs. Reddit comment networks present distinct characteristics in network typological features (e.g., degree, clustering, and density)? Does Moltbook comment network follow typical social network patterns like degree power-law, assortative mixing, and bilateral connection?
@@ 2. Methodology
Moltbook dataset. We construct a Moltbook dataset through Moltbook’s official public API. Specifically, wse first retrieve all posts publicly listed in the Live Posts panel by iteratively calling the post sorting API in chronicle ranking with the offset parameter gradually increased by 100, which gets stopped until the no new posts return. [^1] https://www.moltbook.com/api/v1/posts/?limit=100&sort=new&offset={OFFSET} Afterwards, we retrieve each post’s comments through the post API with its id. [^2] https://www.moltbook.com/api/v1/posts/{POST_ID}/comment In total, we collect 420,259 posts, 2,410,161 comments, and 139,981 replies to comments, posted within the range from $27^{th}$, January, 2026 to $10^{th}$, February, 2026. Figure 1 illlustrates the data schema of Moltbook data.
@@
Inducing comment networks. We construct two directed comment networks for Moltbook and Reddit. A directed link is established from a ”commenter” to a ”recipient” if the former replies to the latter’s post or comment. The resulting Moltbook comment network comprises 39,557 nodes and 697,688 edges, while the Reddit comment network is significantly larger, containing 7,854,970 nodes and 51,850,230 directed edges. Our Moltbook network dataset can be accessed through: https://github.com/James-ZYM/moltbook-typology-analysis
@@ 5. Conclusion
This paper has presented a large-scale analysis of the typology of a online social network populated by AI agents, using Moltbook as a case study. By comparing the comment networks in Reddit and Moltbook, we identify distinct differences in network structural patterns and the efficacy of network building by content posting. In summary, Moltbook network demonstrates a strong hub-and-spoke pattern with a majority of connections are one-way flow. Content posting on Moltbook acts as a catalyst that can initiate link formation consistently, but soon reaches a structural dead end due to the rapid exhaustion of algorithmic attention showing a flash-in-the-pan pattern. These results suggest that, rather than building a social network where connections are persistent and evolve over time, Moltbook functions as a high-speed distribution system characterized by instantaneous but fleeting connection formation. Therefore, Moltbook currently is operated more like an information broadcast network than a robust social community.

### v2

@@ 1. Introduction
To bridge this gap, we conduct the first large-scale comparative analysis on network topology for AI-driven vs. human-drive online social networks. Using Moltbook, a Reddit-like platform operated by AI agents, we collect data from 39K+ AI agents, covering 420K+ posts and 2.5M+ comments generated by agents. For comparison, we gather a parallel dataset from Reddit, comprising 9.3M+ posts and 67M+ comments by 7.8M+ users. Utilizing these datasets, we introduce two comment network mapping commenting interaction for Moltbook agents and Reddit users respectively. We answer following research questions:
  • RQ1: Do Moltbook vs. Reddit comment networks present distinct characteristics in network topological features (e.g., degree, clustering, and density)? Does Moltbook comment network follow typical social network patterns like degree power-law, assortative mixing, and bilateral connection?
@@ 2. Methodology
Moltbook dataset. We construct a Moltbook dataset through Moltbook’s official public API. Specifically, wse first retrieve all posts publicly listed in the Live Posts panel by iteratively calling the post sorting API in chronicle ranking with the offset parameter gradually increased by 100, which gets stopped until the no new posts return. [^1] https://www.moltbook.com/api/v1/posts/?limit=100&sort=new&offset={OFFSET} Afterwards, we retrieve each post’s comments through the post API with its id. [^2] https://www.moltbook.com/api/v1/posts/{POST_ID}/comment In total, we collect 420,259 posts, 2,563,222 comments, posted within the range from $27^{th}$, January, 2026 to $10^{th}$, February, 2026. Figure 1 illlustrates the data schema of Moltbook data.
@@
Inducing comment networks. We construct two directed comment networks for Moltbook and Reddit. A directed link is established from a ”commenter” to a ”recipient” if the former replies to the latter’s post or comment. The resulting Moltbook comment network comprises 39,557 nodes and 697,688 edges, while the Reddit comment network is significantly larger, containing 7,854,970 nodes and 51,850,230 directed edges. Our Moltbook network dataset can be accessed through: https://github.com/James-ZYM/moltbook-topology-analysis
@@ 5. Conclusion
This paper has presented a large-scale analysis of the topology of a online social network populated by AI agents, using Moltbook as a case study. By comparing the comment networks in Reddit and Moltbook, we identify distinct differences in network structural patterns and the efficacy of network building by content posting. In summary, Moltbook network demonstrates a strong hub-and-spoke pattern with a majority of connections are one-way flow. Content posting on Moltbook acts as a catalyst that can initiate link formation consistently, but soon reaches a structural dead end due to the rapid exhaustion of algorithmic attention showing a flash-in-the-pan pattern. These results suggest that, rather than building a social network where connections are persistent and evolve over time, Moltbook functions as a high-speed distribution system characterized by instantaneous but fleeting connection formation. Therefore, Moltbook currently is operated more like an information broadcast network than a robust social community.